# How does the entanglement entropy of a many-body quantum system change after a single measurement?

### v2

@@ IV The QJ measurement protocol
where $\tau$ is the time between quantum jumps and $\Delta S_{nH}=S_{A}(t+\tau)-S_{A}(t)$ computed from Eq. (7). Therefore, $\delta S_{nH}$ captures the system intrinsic capacity to regenerate entanglement between measurements. Since we are interested in the late time dynamics of entanglement for which the average EE has reached the saturation value and therefore does not have a net growth in time, we expect [64] that, on average, the generation of entanglement due to the non-Hermitian evolution will compensate its destruction due to measurements $\overline{\delta S_{nH}}\approx\overline{\Delta S_{qj}}/\bar{\tau}$, where $\overline{\leavevmode\nobreak\ \bullet\leavevmode\nobreak\ }$ stands for the average over quantum trajectories. Results depicted in the left plot of Fig. 8 confirm that this is the case for all considered monitoring strengths. We stress that this cancellation is a property of the average only. We shall see that the distribution of $\delta S_{nH}$ and other properties are qualitatively different from those due to quantum jumps $\Delta S_{qj}$.
@@
Figure 8: The average change in EE after a measurement as a function of the monitoring rate $\gamma$ using the 8 QJ protocol (left) and the 8 PM protocol (right). This is compared with the average speed of entanglement gain between measurements in the QJ protocol $\delta S_{nH}$ (left) and the PM protocol ($\delta S_{un}$) (right). As is expected, the creation of entanglement is compensated by the loss due to the measurement. The system size is $L=512$.
In that region, it is expected that the regeneration rate of EE dominates over its measurement-induced destruction rate governed by the quantum jumps. More specifically, in the weak monitoring limit, the change of EE due to each measurement $\overline{\Delta S_{qj}}$ shows a weak dependence on $\gamma$, while $\bar{\tau}\propto 1/\gamma$, resulting in $\overline{\Delta S_{qj}}/\bar{\tau}\propto\gamma$. To compensate for the destruction of entanglement in this region $\overline{\delta S_{nH}}\propto\gamma$, which is consistent with the results of Fig. 8. When $\gamma$ is sufficiently large, the typical length $\xi$ becomes on the order of or smaller than the lattice constant. In this regime, the spatial extent for entanglement regeneration is suppressed, and the destruction of entanglement by frequent measurements outpaces regeneration. Consequently, $\overline{\delta S_{nH}}$ decreases as $\gamma\gtrsim 1$ increases while the effects of single local measurements $\overline{\Delta S_{qj}}\rightarrow 0$. The observation of a local maximum at $\gamma\sim 3$ in Fig. 8 confirms this picture.
Figure 9: The probability distribution of the change of EE $\delta S_{nH}$ Eq. (10), computed using Eq. (7), between measurements due to the non-hermitian evolution in the QJ protocol for different system sizes and times longer than the saturation time. The monitoring rates are $\gamma=0.1,0.5,1.5$ and $3.0$ from 9 to 9. The monitoring rate is the same as in Fig. 4.
@@
We now address the site dependence of $\delta S_{nH}$, which reveals the regeneration of entanglement after measuring the selected sites. An important difference with respect to the quantum jump contribution $\Delta S_{qj}$ is that, see Fig. 10, $\delta S_{nH}$ is identical for all sites. The stark difference between boundary and bulk sites observed for the change in EE induced by quantum jumps is not present at all in the non-Hermitian contribution to the EE. The absence of any site dependence in $\delta S_{nH}$ indicates that earlier measurements $\Delta S_{qj}$ do not affect the later non-Hermitian evolution $\delta S_{nH}$. $\delta S_{nH}$ is an intrinsic property of the system that only depends on the typical length $\xi$ of the entanglement between the two subsystems, provided that the system is in the area law phase which we believe [41] is always the case for sufficiently large $L$. This qualitative difference between the different contributions to the EE change is a further illustration of the importance, and relevance, of studying local distribution functions for a full characterization of the entanglement dynamics in a many-body quantum system.
@@
Figure 12: The mutual information $I(d_{r})$ Eq. (11) between the subsystem $A$ and a site $M_{r}$ in the other subsystem as a function of the distance $d_{r}$ for the QJ protocol. The pink dashed line is a power-law fitting and the red dashed line is an exponential fitting. The system size is $L=512$.
@@
For weak monitoring, the power-law decay in the average mutual information is a signature of long-range entanglement and therefore of a EE (volume law) that scales with system size. However, previous results in the paper considering the full distribution function, instead of the average, point to a more nuanced picture. Even for weak monitoring, the distribution of $\Delta S_{qj}$ has a peak at zero which increases with the system size. The change in EE due to a quantum jump is specially pronounced in the boundary points, for the rest of sites, it has a relatively narrow distribution around zero. Therefore, even a weak monitoring strength has a profound impact in on the entanglement dynamics that is not captured by the average of the mutual information,
@@ V The projective measurement protocol
Figure 13: The probability distribution $P(\Delta S_{pm})$ of the entanglement entropy change $\Delta S_{pm}$ Eq. (12), computed from Eq. (7) after saturation of the EE, due to a projective measurement of the occupation number (PM protocol). The monitoring rates are $\gamma=0.1,0.5,1.5$ and $3.0$ from 13 to 13.
@@
Figure 15: The density-map of $\delta S_{un}$ Eq. ( 12) for the PM protocol as a function of the occupation number right before the measurement. The monitoring rates are $\gamma=0.1$ (left), and $\gamma=0.5$ (right).
@@ VI Discussion and Conclusions
An interesting question to ask is whether these distributions could provide early signals of a MIPT. As is mentioned in the introduction, at least in the PM [41], and likely in the rest of protocols as well, the system is in the area law phase in the limit of $L\to\infty$ for any $\gamma>0$. For finite $L$, one could still observe a crossover between the area-law a volume-law, or critical, phases as the monitoring rate is decreased. Therefore, the differences observed in the distributions by tuning the monitoring rate suggest that each phase may have distinct features. Not only that, density-maps of the change of the EE before and after a measurement show that even for relatively small $\gamma$ there is a crossover between a peak at half filling associated with the volume law, or critical, to a peak at $n_{i}=1$ (QJ), or $n_{i}=0,1$ (PM), signaling the building up of quantum Zeno effect and therefore of the area law phase. It would be necessary a careful, and numerically costly, finite size scaling analysis to determine whether a transition occurs at a finite $\gamma$ in order to confirm the analytic results of Ref. [41] but our findings provide evidence that, unlike the average, the full distribution function shows the building up of the area-law phase even for relatively small values of the monitoring strength and the system size.
By contrast, for the QSD protocol, we do not find any sign of the area law phase for weak monitoring. However, we have provided arguments that, for this protocol, the distribution of changes is not relevant for the identification of the transition at least in the weak monitoring limit.
@@ A.1 Quantum state diffusion (QSD)
$d({\mathrm{l}n}(|\psi(t)\rangle))=\frac{d|\psi(t)\rangle}{|\psi(t)\rangle}-\frac{\left(d|\psi(t)\rangle\right)^{2}}{2|\psi(t)\rangle^{2}}=\frac{d|\psi(t)\rangle}{|\psi(t)\rangle}-\sum_{i=1}^{L}\left(\frac{\gamma}{2}\left[\hat{n}_{i}-\left\langle\hat{n}_{i}\right\rangle_{t}\right]^{2}\mathrm{\leavevmode\nobreak\ d}t\right)$ (15)
@@
5. Repeat: $2-4$ are repeated for each time step replacing $1$ by the state at the previous time step until the system reaches the steady state characterized by the saturation of the EE $\overline{S(\ell=L/2,t)}$ introduced in Sec II.
@@
6:     Calculate the local occupation number $\langle\hat{n}_{i}\rangle=\sum_{j=1}^{N}U_{ij}U_{ji}^{\dagger}$.
@@ A.2 Quantum jump (QJ)
where $dw_{j}(t)$ is the site-independent Poisson process. Numerically, $dw_{j}(t)$ are discrete random variables that takes only 00 or $1$. Its ensemble average satisfies $\langle dw_{j}(t)\rangle=\gamma dt$, where $\gamma$ is the strength of the measurement that controls the rate at which the measurements are performed.
@@ A.3 Projective Measurements (PM)
The possible outcomes of this measurement are either $1$ or 00, reflecting the fermionic nature of the particles. Then, we can get the density-density correlation matrix $D_{ij}={\rm Tr}(\hat{\rho}\hat{n}_{i}\hat{n}_{j})-{\rm Tr}(\hat{\rho}\hat{n}_{i}){\rm Tr}(\hat{\rho}\hat{n}_{j})$ with $\hat{\rho}$ the density matrix.

### v3

@@ IV The quantum jump measurement protocol
where $\tau$ is the time between quantum jumps and $\Delta S_{nH}=S_{A}(t+\tau)-S_{A}(t)$ computed from Eq. (7). Therefore, $\delta S_{nH}$ captures the system intrinsic capacity to regenerate entanglement between measurements. Since we are interested in the late time dynamics of entanglement for which the average EE has reached the saturation value and therefore does not have a net growth in time, we expect [68] that, on average, the generation of entanglement due to the non-Hermitian evolution will compensate its destruction due to measurements $\overline{\delta S_{nH}}\approx\overline{\Delta S_{qj}}/\bar{\tau}$, where $\overline{\penalty 10000\ \bullet\penalty 10000\ }$ stands for the average over quantum trajectories. Results depicted in the left plot of Fig. 9 confirm that this is the case for all considered monitoring strengths. We stress that this cancellation is a property of the average only. We shall see that the distribution of $\delta S_{nH}$ and other properties are qualitatively different from those due to quantum jumps $\Delta S_{qj}$.
@@
Figure 9: The average change in EE after a measurement as a function of the monitoring rate $\gamma$ using the 9 QJ protocol (left) and the 9 PM protocol (right). This is compared with the average speed of entanglement gain between measurements in the QJ protocol $\delta S_{nH}$ (left) and the PM protocol ($\delta S_{un}$) (right). As is expected, the creation of entanglement is compensated by the loss due to the measurement. The system size is $L=512$.
In that region, it is expected that the regeneration rate of EE dominates over its measurement-induced destruction rate governed by the quantum jumps. More specifically, in the weak monitoring limit, the change of EE due to each measurement $\overline{\Delta S_{qj}}$ shows a weak dependence on $\gamma$, while $\bar{\tau}\propto 1/\gamma$, and therefore $\overline{\Delta S_{qj}}/\bar{\tau}\propto\gamma$. This destruction of entanglement is compensated by $\overline{\delta S_{nH}}\propto\gamma$, which is consistent with the results of Fig. 9. When $\gamma$ is sufficiently large, the typical length $\xi$ becomes on the order of or smaller than the lattice constant. In this regime, the spatial extent for entanglement regeneration is suppressed, and the destruction of entanglement by frequent measurements outpaces regeneration. Consequently, $\overline{\delta S_{nH}}$ decreases as $\gamma\gtrsim 1$ increases while the effects of single local measurements $\overline{\Delta S_{qj}}\rightarrow 0$. The observation of a local maximum at $\gamma\sim 3$ in Fig. 9 confirms this picture.
Figure 10: The probability distribution of the change of EE $\delta S_{nH}$ Eq. (10), computed using Eq. (7), between measurements due to the non-hermitian evolution in the QJ protocol for different system sizes and times longer than the saturation time. The monitoring rates are $\gamma=0.1,0.5,1.5$ and $3.0$ from 10 to 10. The monitoring rate is the same as in Fig. 4.
@@
We now address the site dependence of $\delta S_{nH}$, which reveals the regeneration of entanglement after measuring the selected sites. An important difference with respect to the quantum jump contribution $\Delta S_{qj}$ is that, see Fig. 11, $\delta S_{nH}$ is identical for all sites. The stark difference between boundary and bulk sites observed for the change in EE induced by quantum jumps is not present at all in the non-Hermitian contribution to the EE. The absence of any site dependence in $\delta S_{nH}$ indicates that earlier measurements $\Delta S_{qj}$ do not affect the later non-Hermitian evolution $\delta S_{nH}$.
$\delta S_{nH}$ is an intrinsic property of the system that depends solely on a typical length scale governing the spread of entanglement between two consecutive measurements near the boundary of the two subsystems. This length scale captures the short-time entanglement dynamics and is different from the finite correlation length [37] characterizing the area-law phase of the dynamical equilibrium state.
This qualitative difference between the different contributions to the EE change is a further illustration of the importance, and relevance, of studying local distribution functions for a full characterization of the entanglement dynamics in a many-body quantum system.
@@
Figure 13: The mutual information $I(d_{r})$ Eq. (11) between the subsystem AA and a site $M_{r}$ in the other subsystem as a function of the distance $d_{r}$ for the QJ protocol. The pink dashed line is a power-law fitting and the red dashed line is an exponential fitting. The system size is $L=512$.
@@
For weak monitoring, the power-law decay in the average mutual information is a signature of long-range entanglement and therefore of a EE that scales with the logarithmic of the system size. This is expected because [37] for small $\gamma$ the correlation length is too large to observe the area law phase in the range of sizes we can reach numerically. However, previous results in the paper considering the full distribution function, instead of the average, point to a more nuanced picture. Even for weak monitoring, the distribution of $\Delta S_{qj}$ has a peak at zero which increases with the system size. The change in EE due to a quantum jump is specially pronounced in the boundary points, for the rest of sites, it has a relatively narrow distribution around zero. The decay of the mutual information $I(d_{r})$ further distinguishes the two regimes [37]. In the logarithmic growth regime, where $I(d_{r})$ decays as a power law due to finite size effects, $P(\Delta S_{qj})$ exhibits a power-law divergent peak around $\Delta S_{qj}=0$. This explains that, increasing with system size, a soft peak occurs in the weak monitoring limit. In contrast, in the strong monitoring limit, where $I(d_{r})$ decays exponentially, $P(\Delta S_{qj})$ collapses into a $\delta$-function peak at $\Delta S_{qj}=0$. Therefore, even a weak monitoring strength has a profound impact on the entanglement dynamics that is not captured by the average of the mutual information,
@@ V The projective measurement protocol
Figure 14: The probability distribution $P(\Delta S_{pm})$ of the entanglement entropy change $\Delta S_{pm}$ Eq. (12), computed from Eq. (7) after saturation of the EE, due to a projective measurement of the occupation number (PM protocol). For convenience, the distribution $P(\Delta S_{pm})$ is multiplied by the system size $L$. The monitoring rates are $\gamma=0.1,0.5,1.5$ and $3.0$ from 14 to 14.
@@
Figure 16: The density-map of $\delta S_{un}$ Eq. ( 12) for the PM protocol as a function of the occupation number right before the measurement. The monitoring rates are $\gamma=0.1$ (left), and $\gamma=3.0$ (right).
@@ VI Discussion and Conclusions
An interesting question to ask is whether these distributions could provide early signals of area-law phase. As is mentioned in the introduction, at least in the PM [37], and likely in the rest of protocols as well, the system is in the area law phase in the limit of $L\to\infty$ for any $\gamma>0$. For a fixed $\gamma$, there exists a crossover between a phase where the scaling of the EE is logarithm with the system size, usually termed critical, and a area-law phase as $L$ is increased. The typical length that controls this crossover [37] grows exponentially with $1/\gamma$ so deviations from the phase characterized by a logarithmic growth of the EE are small for $\gamma\ll 1$. However, our results for the density-maps of the change of the EE before and after a measurement show that even for relatively small $\gamma$ there is a crossover between a peak at half filling associated with the logarithmic growth with system size to a peak at $n_{i}=1$ (QJ), or $n_{i}=0,1$ (PM), signaling the building up of the quantum Zeno effect. A demonstration, based on a finite size scaling analysis, that the observed Zeno-like features in the distribution function develop into an area law for the EE for any $\gamma>0$ is numerically costly and beyond the scope of the paper. A finite size scaling analysis of just the average [37], not the full distribution, is enough to confirm the absence of a MIPT. However, our findings provide evidence that, unlike the average, the distribution function shows changes, even for small values of $\gamma$ and not very large system sizes, which suggests that monitoring is already strongly influencing the entanglement dynamics. By contrast, for the QSD protocol, we do not find any sign of the area law phase for weak monitoring. However, we have provided arguments that, for this protocol, the distribution of changes is not relevant for the identification of the transition at least in the weak monitoring limit.
@@ A.1 Quantum state diffusion (QSD)
$d({\mathrm{l}n}(|\psi(t)\rangle))=\frac{d|\psi(t)\rangle}{|\psi(t)\rangle}-\frac{\left(d|\psi(t)\rangle\right)^{2}}{2|\psi(t)\rangle^{2}}=\frac{d|\psi(t)\rangle}{|\psi(t)\rangle}-\sum_{i=1}^{L}\left(\frac{\gamma}{2}\left[\hat{n}_{i}-\left\langle\hat{n}_{i}\right\rangle_{t}\right]^{2}\mathrm{\penalty 10000\ d}t\right)$ (15)
@@
5. Repeat: 242-4 are repeated for each time step replacing $1$ by the state at the previous time step until the system reaches the steady state characterized by the saturation of the EE $\overline{S(\ell=L/2,t)}$ introduced in Sec II.
@@
6:  Calculate the local occupation number $\langle\hat{n}_{i}\rangle=\sum_{j=1}^{N}U_{ij}U^{*}_{ij}$.
@@ A.2 Quantum jump (QJ)
where $dw_{j}(t)$ is the site-independent Poisson process. Numerically, $dw_{j}(t)$ are discrete random variables that takes only 0 or $1$. Its ensemble average satisfies $\langle dw_{j}(t)\rangle=\gamma dt$, where $\gamma$ is the strength of the measurement that controls the rate at which the measurements are performed.
@@ A.3 Projective Measurements (PM)
The possible outcomes of this measurement are either $1$ or 0, reflecting the fermionic nature of the particles. Then, we can get the density-density correlation matrix $D_{ij}={\rm Tr}(\hat{\rho}\hat{n}_{i}\hat{n}_{j})-{\rm Tr}(\hat{\rho}\hat{n}_{i}){\rm Tr}(\hat{\rho}\hat{n}_{j})$ with $\hat{\rho}$ the density matrix.